What is the free cash flow of a firm with revenues of $386 million, operating profit margin of 31%, tax rate of 34%, depreciation and amortization expense of $21 million, capital expenditures of $39 million, acquisition expenses of $5 million and change in net working capital of $19 million? Answer in millions, rounded to one decimal place (e.g., $245.63 = 245.6).
Answer #1

Calculation of free cashflow:

Free cashflow= operating profit- tax+depreciation-Capital expenditure- change in net working capital

Operating profit= 386*0.31= $119.66 million

Tax= 119.66*0.34= 40.68

Free cashflow= 119.66-40.68+21-39-19= 41.98

Free cashflow= 42 million